OpenStudy (anonymous):

how do you do. (xy^6)^-2

OpenStudy (jdoe0001):

\(\Large a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (anonymous):

would i set it up (xy^6)/2 then?

OpenStudy (anonymous):

indices of variables inside brackets are multiplied by the indice outside the bracket. As such this equals (x^(1*-2)) * (y^(6*-2)) = x^-2 * y^-12 Using rule posted above this equals 1/(x^2*y^12) If using rule posted by jdoe001 above from the start.... 1/(x * y^6)^2 = 1 divided by { (x^(1*2)) * (y^(6*2)) ] Note x = x^1, they are the same thing.

OpenStudy (anonymous):

i see thank you